# Funding AI for Good: A Call for Meaningful Engagement

### v2

@@ Abstract.
Artificial Intelligence for Social Good (AI4SG) is a growing area that explores AI’s potential to address social issues, such as public health. Yet prior work has shown limited evidence of its tangible benefits for intended communities, and projects frequently face inadequate community engagement and sustainability challenges. While existing HCI literature on AI4SG initiatives primarily focuses on the mechanisms of funded projects and their outcomes, much less attention has been given to the funding agenda and rhetoric that influences downstream approaches. Through a thematic analysis of 35 funding documents—representing about \$410 million USD in total investments, we reveal dissonances between AI4SG’s stated intentions for positive social impact and the techno-centric approaches that some funding agendas promoted, while also identifying funding documents that scaffolded community-collaborative approaches for applicants. Drawing on our findings, we offer recommendations for funders to embed approaches that balance both contextual understanding and technical capacities in future funding call designs. We further discuss how the HCI community can positively shape AI4SG funding design processes.
copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXXprice: 15.00isbn: XXXXX
@@ 1. Introduction
Artificial Intelligence (AI) for social good (AI4SG) initiatives aim to create long-lasting positive impact in social domains such as public health, by leveraging AI technologies like machine learning (Shi et al., 2020; Li et al., 2021; Perrault et al., 2020; Rolnick et al., 2022; Tomašev et al., 2020). Yet prior work finds limited evidence of its tangible benefits for intended communities, and projects frequently face deployment and sustainability challenges in real-world contexts (Vinuesa et al., 2020; Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Tate Ryan-Mosley, 2023). HCI scholarship provides insights into the socio-technical mechanisms underpinning these challenges, and recent work points to the significant influence of funding agendas—the priorities, goals, and approaches set by funders—in shaping the technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). In AI4SG specifically,  Lin et al. (2024) show from community organizations’ perspectives that funding agendas heavily influence how partnerships are formed, which problem spaces are prioritized, and what kinds of solutions are pursued during the early stages of projects. Prior work also notes that funding often incentivizes mission-driven organizations to adopt data-driven innovations, sometimes at the cost of their original social missions (Bopp et al., 2017; Lin et al., 2024). For AI developers and practitioners, short funding timelines can constrain considerations of ethics and unintended consequences (Do et al., 2023; Moitra et al., 2022; Okolo and Lin, 2024). However, existing HCI literature on AI4SG initiatives primarily focuses on funded projects’ approaches and outcomes through the perspectives of project teams (typically consisting of AI technologists and mission-driven organizations); much less attention is given to the design of funding agendas themselves, before projects even get selected.
@@
In this work, we begin addressing this gap by investigating AI4SG funding agendas, following up on Laura Nader’s call to “study up” (Nader, 1972) systems and actors that exert normative influence. In doing so, we aim to shift attention from assigning blame to AI4SG project teams, and towards recognizing the broader power structures shaping their approaches. We turn to public funding documents—textual instruments that communicate funding agendas, including calls for proposals and grant announcements. Prior literature has established that funding documents are not neutral instruments: they construct priorities (Palmero et al., 2025), shape applicant behaviors (Peng et al., 2024), and structure downstream research activities (Widder et al., 2024; Smith et al., 2023). In AI4SG, by defining who is eligible, what constitutes a “problem,” how these problems should be addressed, and what “success” means, funding documents construct discourses about AI and social good, signaling what approaches are fundable or imaginable. Although there is complexity in how written funding documents influence what actually gets funded, examining them reveals how priorities, values, and power relations are embedded even before any project begins. Funding documents can reinforce dominant AI imaginaries—“collective visions, beliefs, symbols, and expectations that individuals and communities hold about AI and its potential outcomes” (Zhong et al., 2025)—as seen in public texts portraying AI as “inevitable and massively disrupting” (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022). Yet they can also guide applicants toward community engagement strategies previously unfamiliar to them.
@@
  • RQ1: How do funding documents frame social good, and what is AI’s envisioned role in advancing social good, as articulated in funding documents?
@@
  • RQ3: How do funding documents characterize how these goals might be achieved?
@@
Since AI4SG builds on earlier “tech for good” efforts (Aula and Bowles, 2023), we ground our analysis on lessons learned from Computer-Supported Cooperative Work (CSCW), Information and Communication Technologies and Development (ICTD), and Human-Computer Interaction for Development (HCI4D). This prior literature has extensively examined different approaches in socially driven technology projects, emphasizing that impactful innovations require both a thorough understanding of the problem and appropriate technological capabilities (Figure 1). From our synthesis of this body of work, a spectrum from techno-centric to balanced approaches emerges, indicating a strong association between where projects fall on this spectrum and project outcomes. A techno-centric approach assumes novel technology alone can solve complex social issues, often producing suboptimal results and strained relationships between researchers and communities (e.g., (Ames, 2019; Karusala and Kumar, 2017; Walsham, 2017; Dada, 2006; Brown and Mickelson, 2019)). In contrast, a more balanced approach, grounded in a solid understanding of the community’s challenges and contexts through community-collaborative approaches (Cooper and Zafiroglu, 2024), tend to result in more community buy-in and sustainability (e.g., (Toyama, 2015; Gandhi et al., 2007; Burrell and Toyama, 2009)). Using this spectrum, our study analyzes how current AI4SG funding rhetoric is oriented towards a more techno-centric or a more balanced approach.
Funding documents in our corpus overwhelmingly mobilized AI against large-scale social problems that cannot be fully addressed with any single intervention (notably public health and development). Yet several funding documents framed AI’s potential as deterministically beneficial and positioned AI interventions as solutions to these social problems while skirting contextual and stakeholder specificity. Some other documents scoped interventions to more well-defined sub-problems and explicitly acknowledged both AI’s benefits and risks. Expected outcomes of funding documents in our dataset frequently centered on technical deliverables, while a subset adopted broader success metrics attentive to benefits to the communities involved. Furthermore, although interdisciplinary collaboration was widely encouraged, funding documents varied in their emphasis and scaffolding of community engagement and who were considered “experts.” With a few exceptions, expectations for community engagement stopped at a consultation mode towards later stages of projects (e.g., implementation stage), where community stakeholders were expected to provide inputs but not able to shape the broader project scopes.
@@
First, we provide a novel focus on AI4SG funding documents as a discourse instrument rather than neutral texts. Through a reflexive thematic analysis (Braun and Clarke, 2006, 2019), we highlight the disconnects between the positive social impact AI4SG aims for and the techno-centric approaches promoted in some funding documents (Section 5.1). Second, building on the more balanced approaches and gaps identified in our dataset, we propose opportunities for designing future funding calls for funders (including supporting relationship-building, maintenance post-project deployment, and community engagement training for project teams), and outline future research directions for HCI researchers (Section 5.2). Finally, we call for meaningful community engagement in the funding design process for funders, and specify the roles that HCI practitioners and researchers can play in shaping AI4SG funding agendas and design processes (Section 5.3).
@@ 2. Related Work
Prior HCI research on AI4SG has highlighted both the objective of existing funded projects to produce positive social impact and the challenges they face in doing so (Section 2.1). Scholars have called on AI4SG to draw on lessons from previous “tech for good” projects, which emphasize that impactful innovations require a balance of appropriate technical capacity and deep contextual understanding (Section 2.2). Given the central role of funding agendas in shaping the AI4SG ecosystem, we draw on emerging scholarship examining grant-making and private tech companies’ investment in AI4SG (Section 2.3). Furthermore, prior research in HCI and innovation studies has enacted document analysis to examine how AI is imagined in public texts and established the role of funding documents in shaping research priorities and collaboration structures (Section 2.4). However, HCI has yet to systematically analyze funding documents that define and construct AI4SG initiatives. This study begins to address this gap by analyzing AI4SG funding documents through the lens of existing literature on “tech for good.” While prior work points out funded projects’ bias towards emphasizing technical capacity, our work contributes an analysis of how this bias is instantiated across some funding documents, while also identifying more balanced approaches that emphasize contextual understanding.
@@ 2.1. The Promises and Challenges of AI for Social Good
Artificial Intelligence for Social Good (AI4SG) has emerged as a growing field of use-inspired research and real-world implementations focused on applying AI to complex social issues (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). Its stated goal is to deliver transformative, positive, and lasting impact across many social aspects (Perera, 2024; Cowls, 2021; Tudor et al., 2024). Some notable examples include projects in agriculture that predict crop yields using remote sensing data (You et al., 2017) and provide price forecasting for farmers in India (Ma et al., 2019). While enthusiasm for AI4SG remains high, growing evidence shows that many projects fall short of delivering lasting positive social impact and often end at the publication stage without deployment (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). In high-stakes contexts such as public service screening or criminal justice, AI may be inappropriate altogether (Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Pruss, 2023). Additionally, AI4SG is often shaped by promotional hype (Joyce and Cruz, 2024), magical thinking (Lupetti and Murray-Rust, 2024), and “enchanted determinism” (Campolo and Crawford, 2020), where the fascination with the mysterious capabilities of new technologies captures public attention toward techno-solutionism—where complex, structural challenges are reduced to problems amenable to AI interventions (Morozov, 2013). This could be exacerbated by AI’s transparency issues (especially in the Global South context that AI4SG projects are intended for (Okolo and Lin, 2024)). For instance, studies in India show a strong belief in “AI authority” over public institutions and lowered expectations for accountability (Kapania et al., 2022; Ramesh et al., 2022).
@@
HCI scholars have started to examine the mechanisms behind AI4SG projects. We highlight two key insights below. First, the aspiration to employ AI for social good is not a recent phenomenon; AI4SG should draw lessons learned from prior “tech for good” movements (Aula and Bowles, 2023; Lin et al., 2024). ICTD and HCI4D have long explored the development and use of emerging technologies in marginalized communities, especially in the Global South (Walsham, 2017; Dell and Kumar, 2016; Avgerou, 2010). Second, while past work often focuses on investigating perspectives of project teams—typically AI developers and nonprofit organizations, emerging research highlights funding agendas’ role in shaping interdisciplinary partnerships (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). Through interviews with staff of community organizations participating in AI4SG,  Lin et al. (2024) found that funding agenda influences program goals, proposed solutions, and definitions of success. Focusing on nonprofit organizations, other studies show how funding agendas shape their incentives to adopt emerging tech, and collaboration choices (Bopp et al., 2017; Erete et al., 2016; Goldenfein and Mann, 2023). AI developers and researchers’ approaches to AI ethics and considerations of unintended consequences are also shown to be impacted by funding timeline (Do et al., 2023; Widder et al., 2023; Moitra et al., 2022).
@@
Emphasizing the need to examine how power structures impact technology development in social contexts, scholars have taken up Laura Nader’s (1972) call to “study up” by focusing on systems and actors that shape social systems and exert normative influence (Nader, 1972). Emerging “study up” scholarship in algorithmic fairness (Barabas et al., 2020), public sector AI (Kawakami et al., 2024a), and machine learning data (Miceli et al., 2022) demonstrates that examining those in power can reveal critical leverage points for more accountable AI development. While prior work has highlighted funding’s significant influence on project approaches and outcomes, we still lack a direct understanding of how funding frames AI’s role in addressing social good challenges; there is a need to “study up” AI4SG funding.
@@ 2.2. Lessons from Prior “Tech for Good” Projects: Techno-Centric vs Balanced Approaches
In this section, we synthesize this rich body of work and categorize characteristics of projects into a more techno-centric approach (bias towards technical solutions) and a more balanced approach (emphasizing both technical capacities and contextual understanding). We particularly focus on aspects of motivations, collaborations, longevity and process, and evaluations, because these aspects are closely influenced by funding agendas (Lin et al., 2024).

### v3

@@ Abstract.
Artificial Intelligence for Social Good (AI4SG) is a growing area that explores AI’s potential to address social issues, such as public health. Yet prior work has shown limited evidence of its tangible benefits for intended communities, and projects frequently face real-world deployment and sustainability challenges. While existing HCI literature on AI4SG initiatives primarily focuses on the mechanisms of funded projects and their outcomes, much less attention has been given to the upstream funding agendas that influence project approaches. In this work, we conducted a reflexive thematic analysis of 35 funding documents, representing about \$410 million USD in total investments. We uncovered a spectrum of conceptual framings of AI4SG and the approaches that funding rhetoric promoted: from biasing towards technology capacities (more techno-centric) to emphasizing contextual understanding of the social problems at hand alongside technology capacities (more balanced). Drawing on our findings on how funding documents construct AI4SG, we offer recommendations for funders to embed more balanced approaches in future funding call designs. We further discuss implications for how the HCI community can positively shape AI4SG funding design processes.
AI for Social Good; Funding; Document Analysis; AI imaginaries
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790374isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
This figure shows a spectrum of two approaches in pursuing technology development for social impact. The left panel shows an overemphasis on technical problems (a more techno-centric approach), which could lead to innovations without social impact. Our work found that some AI4SG funding documents led with rhetoric of more techno-centric approaches that are biased towards technological capacities. The right panel shows a balance of emphasis on technical capacities and contextual understanding (a more balanced approach), which could lead to impactful innovations. This work offers recommendations to bring AI4SG funding agendas closer to more balanced approaches that emphasize both technology capacities and contextual understanding.
@@
Artificial Intelligence (AI) for social good (AI4SG) initiatives aim to create long-lasting positive impact in social domains such as public health, by leveraging AI technologies like machine learning (Shi et al., 2020; Li et al., 2021; Perrault et al., 2020; Rolnick et al., 2022; Tomašev et al., 2020). Yet prior work finds limited evidence of tangible benefits of these initiatives for intended communities, and projects frequently face deployment and sustainability challenges in real-world contexts (Vinuesa et al., 2020; Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Tate Ryan-Mosley, 2023). HCI scholarship provides insights into the socio-technical mechanisms underpinning these challenges, and recent work points to the significant influence of funding agendas—the priorities, goals, and approaches set by funders—in shaping the technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). In AI4SG specifically, Lin et al. (2024) foreground community organizations’ perspectives and show that funding agendas heavily influence how partnerships are formed, which problem spaces are prioritized, and what kinds of solutions are pursued during the early stages of projects. Prior work also notes that funding often incentivizes mission-driven organizations to adopt data-driven innovations, sometimes at the cost of their original social missions (Bopp et al., 2017; Lin et al., 2024). For AI developers and practitioners, short funding timelines can constrain considerations of ethics and unintended consequences (Do et al., 2023; Moitra et al., 2022; Okolo and Lin, 2024). However, existing HCI literature on AI4SG initiatives primarily focuses on funded projects’ approaches and outcomes through the perspectives of project teams; much less attention is given to the design of upstream funding agendas themselves, before projects even get selected.
@@
In this work, we begin addressing this gap by investigating AI4SG funding agendas, following up on Laura Nader’s call to “study up” (Nader, 1972) systems and actors that exert normative influence. In doing so, we aim to shift attention from assigning blame to AI4SG project teams (typically consisting of AI technologists and mission-driven organizations), and towards recognizing the broader power structures shaping their approaches. We turn to public funding documents—textual instruments that construct and communicate funding agendas, including calls for proposals and grant announcements. Prior literature has established that funding documents are not neutral instruments: they construct priorities (Palmero et al., 2025), shape applicant behaviors (Peng et al., 2024), and structure downstream research activities (Widder et al., 2024; Smith et al., 2023). In AI4SG, by defining who is eligible, what constitutes a problem, how these problems should be addressed, and what success means, funding documents construct discourses about “AI” and “social good”, signaling what approaches are fundable or imaginable. Although the mechanisms by which written funding documents influence what actually gets funded are complex, examining funding rhetoric reveals how priorities, values, and power relations are embedded before any project begins. Funding documents can reinforce dominant AI imaginaries—“collective visions, beliefs, symbols, and expectations that individuals and communities hold about AI and its potential outcomes” (Zhong et al., 2025)—as seen in public texts portraying AI as “inevitable and massively disrupting” (Bareis and Katzenbach, 2022). Yet they can also guide applicants toward community engagement strategies previously unfamiliar to them.
@@
  • RQ1: How do funding documents frame social good and AI’s role in advancing social good?
@@
  • RQ3: How do funding documents envision how these goals might be achieved?
@@
Since AI4SG builds on earlier “tech for good” efforts (Aula and Bowles, 2023), we ground our analysis in lessons learned from Computer-Supported Cooperative Work (CSCW), Information and Communication Technologies and Development (ICTD), and Human-Computer Interaction for Development (HCI4D). This prior literature has extensively examined different approaches in socially driven technology projects, emphasizing that impactful innovations require both a thorough understanding of the problem and appropriate technological capabilities (Figure 1). From our synthesis of this body of work, a spectrum from techno-centric to balanced approaches emerges, indicating a strong association between where projects fall on this spectrum and project outcomes. A techno-centric approach assumes novel technology alone can solve complex social issues, often producing suboptimal results and strained relationships between researchers and communities (e.g., (Ames, 2019; Karusala and Kumar, 2017; Walsham, 2017; Dada, 2006; Brown and Mickelson, 2019)). In contrast, a more balanced approach, grounded in a solid understanding of the community’s challenges and contexts through community-collaborative approaches (Cooper and Zafiroglu, 2024), tend to result in more community buy-in and sustainability (e.g., (Toyama, 2015; Gandhi et al., 2007; Burrell and Toyama, 2009)). Using this spectrum, we examine how current AI4SG funding rhetoric is oriented towards a more techno-centric or a more balanced approach.
Funding documents in our corpus overwhelmingly mobilized AI against large-scale social problems that cannot be fully addressed with any single intervention (notably public health and development). Yet several funding documents framed AI’s potential as deterministically beneficial and positioned AI interventions as solutions to these social problems while skirting contextual and stakeholder specificity. Some other documents scoped interventions to more well-defined sub-problems and explicitly acknowledged both AI’s benefits and risks. Expected outcomes expressed in the funding documents in our dataset frequently centered on technical deliverables, while a subset adopted broader success metrics attentive to benefits to the communities involved. Furthermore, although interdisciplinary collaboration was widely encouraged, funding documents varied in their emphasis and scaffolding of community engagement and who were considered “experts.” With a few exceptions, expectations for community engagement stopped at a consultation mode towards later stages of projects (e.g., implementation stage), where community stakeholders were expected to provide inputs but not be able to shape the broader project scopes.
@@
  • First, we provide a novel focus on AI4SG funding documents as a discourse instrument rather than neutral texts. Through a reflexive thematic analysis (Braun and Clarke, 2006, 2019), we highlight the disconnects between the positive social impact AI4SG aims for and the techno-centric approaches promoted in some funding documents (Section 5.1).
  • Second, building on the more balanced approaches and gaps identified in our dataset, we propose opportunities for designing future funding calls for funders (including supporting relationship-building, maintenance post-project deployment, and community engagement training for project teams), and outline future research directions for HCI researchers (Section 5.2).
  • Finally, we call for meaningful community engagement in the funding design process for funders, and specify the roles that HCI practitioners and researchers can play in shaping AI4SG funding agendas and design processes (Section 5.3).
@@ 2. Related Work
We draw on prior HCI research on AI4SG, which has highlighted the objectives of existing funded projects to produce positive social impact, the challenges they face in achieving them, and the central role of funding agendas in shaping the AI4SG ecosystem (Section 2.1). Scholars have called on AI4SG to draw on lessons from previous “tech for good” projects, which emphasize that impactful innovations require a balance of appropriate technical capacity and deep contextual understanding (Section 2.2). We then draw on emerging scholarship examining grant-making and private tech companies’ investment in AI4SG (Section 2.3). Lastly, we draw on prior research in HCI and innovation studies that use document analysis to examine how AI is imagined in public texts, and emphasize the role of funding documents in shaping research priorities and collaboration structures (Section 2.4).
HCI has yet to systematically analyze how funding documents discursively define and construct AI4SG initiatives and the conditions under which their positive impact might be achieved. This study begins to address this gap by analyzing AI4SG funding documents through the lens of existing literature on “tech for good.” While prior work points out funded projects’ bias towards technical capacity, our work contributes an analysis of how this bias is instantiated across some funding documents, while also identifying more balanced approaches that emphasize contextual understanding.
@@ 2.1. The Promises and Challenges of AI for Social Good
Artificial Intelligence for Social Good (AI4SG) has emerged as a growing field of use-inspired research and real-world implementations focused on applying AI to complex social issues (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). Its stated goal is to deliver transformative, positive, and lasting impact across many social aspects (Perera, 2024; Cowls, 2021; Tudor et al., 2024). Some notable examples include projects in agriculture that predict crop yields using remote sensing data (You et al., 2017) and provide price forecasting for farmers in India (Ma et al., 2019). While enthusiasm for AI4SG remains high, growing evidence shows that many projects fall short of delivering lasting positive social impact and often end at the publication stage without deployment (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). In high-stakes contexts such as public service screening or criminal justice, AI may be inappropriate altogether (Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Pruss, 2023). Additionally, AI4SG is often shaped by promotional hype (Joyce and Cruz, 2024), magical thinking (Lupetti and Murray-Rust, 2024), and “enchanted determinism” (Campolo and Crawford, 2020), where the fascination with the mysterious capabilities of new technologies captures public attention toward AI interventions. This could be exacerbated by AI’s transparency issues (especially in the Global South context that AI4SG projects are intended for (Okolo and Lin, 2024)). For instance, studies in India show a strong belief in “AI authority” over public institutions and lowered expectations for accountability (Kapania et al., 2022; Ramesh et al., 2022).
@@
HCI scholars have started to examine the mechanisms underlying AI4SG projects. We highlight two key insights below. First, the aspiration to employ a powerful emerging technology for social good is not a recent phenomenon; AI4SG should draw lessons learned from prior “tech for good” movements (Aula and Bowles, 2023; Lin et al., 2024). ICTD and HCI4D have long explored the development and use of emerging technologies in marginalized communities, especially in the Global South (Walsham, 2017; Dell and Kumar, 2016; Avgerou, 2010). Second, while past work often focused on investigating perspectives of project teams—typically AI developers and nonprofit organizations, emerging research highlights funding agendas’ role in shaping interdisciplinary partnerships (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). Through interviews with staff from community organizations participating in AI4SG, Lin et al. (2024) found that funding agendas influence program goals, proposed solutions, and definitions of success. Focusing on nonprofit organizations, other studies show how funding agendas shape their incentives to adopt emerging tech, and their collaboration choices (Bopp et al., 2017; Erete et al., 2016; Goldenfein and Mann, 2023). AI developers and researchers’ approaches to AI ethics and considerations of unintended consequences are also shown to be impacted by funding timeline (Do et al., 2023; Widder et al., 2023; Moitra et al., 2022).
@@
Emphasizing the need to examine how power structures impact technology development in social contexts, scholars have taken up Laura Nader’s call to “study up” by focusing on systems and actors that shape social systems and exert normative influence (Nader, 1972). Emerging “study up” scholarship in algorithmic fairness (Barabas et al., 2020), public sector AI (Kawakami et al., 2024a), and machine learning data (Miceli et al., 2022) demonstrates that examining those in power can reveal critical leverage points for more accountable AI development. While prior work has highlighted funding’s significant influence on project approaches and outcomes, we still lack a direct understanding of how funding frames AI’s role in addressing social good challenges; there is a need to “study up” AI4SG funding.
@@ 2.2. Lessons from Prior “Tech for Good” Projects: Techno-Centric vs Balanced Approaches
In this section, we synthesize this rich body of work and categorize project characteristics into a more techno-centric approach (bias towards technical solutions) and a more balanced approach (emphasizing both technical capacities and contextual understanding). We particularly focus on aspects of motivation, collaboration, longevity, process, and evaluation because these are closely influenced by funding agendas (Lin et al., 2024).